water of density 4 kg/m3 and ice of density 2 kg/m3 are mixed together. if their masses are equal then the density of mixture is?

Grade:9

1 Answers

Arun
6 years ago
Dear student
 
If twi loquid lf equal mass and different densities are mixed then
Density = 2 *d1 *d2/(d1 +d2)
 = 16/6 kg/m³
 = 8/3 kg/m³
 
Regards
Arun (askIITians forum expert)